OpenStudy (anonymous):

solve the system by the elimination method 5x+5y=-11 7x-2y=19

OpenStudy (anonymous):

multiply first equation by 7 and second by -5 and the add the resulting equations and tell me what u get.... We will proceed from there

OpenStudy (anonymous):

@purplec pls do not provide the solution otherwise lizzy will not learn how to do this kind of sums on her own

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

@lizzy pls do as i suggested and tell me what u get......

OpenStudy (anonymous):

hey lizzy???????????

OpenStudy (anonymous):

10x+5y=-11 21x-6y=19

OpenStudy (anonymous):

this is wrong, u have to multiply all terms by the number try again..... multiply the first equation one by 7 and second equation by -5

OpenStudy (anonymous):

I thought it was by two and three

OpenStudy (anonymous):

why??? our aim is to make the coefficients of x same (we can even make coefficients of y equal if we want...)

OpenStudy (anonymous):

35x+35=-11 -35x++-10y=19

OpenStudy (anonymous):

someone elese on here told me to multiply first line by 2 and second by 3. is this wrong???

OpenStudy (anonymous):

I am so confused now..OMG

OpenStudy (anonymous):

I have nine problems like this and i do not understand how to do them..can u help me

OpenStudy (anonymous):

see it might have been for some other sum..... by the way, do you hv a SKYPE account???? If yes, v can have voice chat n things will move faster...

OpenStudy (anonymous):

what's that SKYPE account???

OpenStudy (anonymous):

is is a downloadable program which u can download from SKYPE.COM website and then u and i can chat over the net...

OpenStudy (anonymous):

my SKYPE id is hakki.singh

OpenStudy (anonymous):

why we just can't chat here??

OpenStudy (anonymous):

ok here we hv to type, that takes time...

OpenStudy (anonymous):

see for such "elimination method" questions, we have to try to make the coefficients of either x or y equal so that we can eliminate one variable.. Understood...

OpenStudy (anonymous):

5x + 5y = -11 -----------(1) 7x - 2y = 19 -----------(2) so in this case we multiply (1) by 7 and (2) by -5, to get 35x + 35y = -77 ---------(3) -35x + 10y = -95 ---------(4) got it???

OpenStudy (anonymous):

Now if we add (3) and (4) we get 45y = -172 y = -172/45

OpenStudy (anonymous):

now put y = -172/45 in (1) -172 5x + 5*------- = -11 45 172 5x - ----- = -11 9 172 5x = -11 + ------ 9 -99 + 172 5x = ----------- 9 73 x = -------- 45

OpenStudy (anonymous):

so final solution is x = 73/45 and y = -172/45